OpenStudy (clamin):

PLEASE HELP!!! MEDAL!!! Please show me the x and y value of y=-x^2+2x-2

OpenStudy (sepeario):

Okay let's solve for x first. Have you heard of the quadratic formula?

OpenStudy (clamin):

Yes. y=ax^2+bx+c @sepeario

OpenStudy (anonymous):

you can break it up like this :) well, you're looking for to this formula : \[x=\frac{ b \pm \sqrt{b ^{2} -4ac}} {2a }\] now let's solve it :) -x^2+2x-2 : this is your equation, now let's break down a little bit: -x^2 = a=(-1): when it's x, you can put it as one to easy to calculate, plus, it make sense, you can use any other real number but the results it not so beautiful... okay, next: 2x = b -2 = c Now we know all the numbers and letters: \[y= \frac{ 2\pm \sqrt{2^{2}-4(-1)(-2)} }{ 2(-1) }\] \[y=\frac{ 2\pm \sqrt{4-8} }{ -2 }\] Since \[\sqrt{4-8} = \sqrt{-4}\] which doesn't make sense, so it will be false! which now we know that it will be NO SOLUTION! :) hope it help :)

OpenStudy (clamin):

This helps a little but I'm looking for the x and y value which I struggle a lot to find..@leong

OpenStudy (clamin):

@Leong

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

you can't look for x and y when you know that's it false, you can try your formula which is (a+b)^2= (a^2+2ab+b^2), I did that and it give me very ugly numbers.....

OpenStudy (anonymous):

not every equations have numbers or results, since the equations is kinda false itself. Like the one you give me, I can show you if you wanted by using others quadratic formula, but trust me, you will ending up with ugly numbers or false :)

OpenStudy (clamin):

Its for the graph

OpenStudy (freckles):

what do you mean to show you the x and y value?

OpenStudy (freckles):

x and y are variables they vary they don't just take on one value

OpenStudy (sepeario):

i think he/she means solve for x and y.

Nnesha (nnesha):

x and y intercepts ? @clamin ?

Nnesha (nnesha):

if yes then use \(\huge\color{reD}{\rm b^2-4ac}\) `Discriminant` if ` b^2-4ac > 0` then there are 2 real zeros if ` b^2-4ac = 0` then there is one real root if ` b^2-4ac < 0` then you will get two complex roots (no -x-intercept) to find y-intercept substitute x for 0 solve for y

OpenStudy (phi):

If you want some (x,y) pairs so that you can graph the equation \[ y=−x^2+2x−2 \] I would start by picking some small x values (they are easier to figure out) for example, when x is 0 we find y by doing this: \[ y=−(0)^2+2⋅0−2 \] hopefully you know that simplifies to y=-2 so (0,-2) is one point

OpenStudy (phi):

let's try x=-1: \( y=−(−1)^2+2⋅−1−2\) notice that when we replaced x with -1 in -x^2 we put in parens -(-1)^2 - (-1)^2 means -1* (-1)^2 or -1* (-1*-1) or -1*1 and finally -1 we get y=−1−2−2=−5 so (-1,-5) is another point

OpenStudy (phi):

now try x=1 y= -(1)^2 +2*1 -2 y= -(1*1) + 2 -2 y= -1 so (1,-1) is a point try x=2 y= -(2*2) + 2*2 -2 = -4+4-2 = -2 (2,-2) x=3 y= -(3*3) +2*3 -2 = -9 +6-2 = -5 (3,-5)

OpenStudy (clamin):

@phi how come the 4 is negative?

OpenStudy (clamin):

@nnesha @freckles @sepeario @leong

OpenStudy (phi):

which 4 ?

OpenStudy (clamin):